OpenStudy (anonymous):

Find the vertical asymptotes, if any, for the following function. Please show all of your work. f(x) = x+11/x^2-16x

OpenStudy (anonymous):

x^2 - 16x = 0

jimthompson5910 (jim_thompson5910):

the denominator cannot be zero if it were, then x^2-16x = 0 x(x-16) = 0 x=0 or x-16=0 x=0 or x=16 So the values x=0 or x=16 cause a division by zero error So the vertical asymptotes are x=0 and x=16

OpenStudy (anonymous):

james I have a question..after your done with this one..can you come back to my question? thank you

OpenStudy (anonymous):

never mind..thank you for helping me with those questions

OpenStudy (anonymous):

i can help you. which question do i go bck to?

OpenStudy (anonymous):

Thank you Jamesm!!